OpenStudy (anonymous):

if tanh(x) =5/13 , find the value of the hyperbolic function of sech(x)

OpenStudy (anonymous):

\[\tanh(x)=\sinh(x)/\cosh(x)=5/13\] \[sech(x)=1/\cosh(x)=1/13\]